OpenStudy (anonymous):

f(x) = 7x + 7, g(x) = 6x2 Find (fg)(x).

OpenStudy (anonymous):

(fg)(x) = f(g(x)) = 7(6x^2) + 7 = 7(6x^2 +1)

OpenStudy (anonymous):

thanks!